# Segmented Operations using Matrix Multiplications

### v1

@@ 2.1 Model definitions: Vector-RAM, TCU, and circuits
  • Threshold circuits are boolean circuits which also include so-called threshold or majority gates, i.e., gates which receive as input $k$ boolean variables $x_{1},\ldots,x_{k}$ and return 1 if and only if $\sum_{i=1}^{k}w_{i}x_{i}\geq t$, for some fixed integers $t,w_{1},\ldots,w_{k}$.
For $i\in\mathbb{N}$, NC[i], AC[i], and TC[i] are the classes of boolean circuits with fan-in two, unbounded fan-in, and threshold circuits, respectively, all having $n^{O(1)}$ size and $\mathcal{O}(\log^{i}(n))$ depth, where $n$ is the input size. It is known that NC[0$\subset$ AC[0] $\subset$ TC[0] $\subseteq$ NC[1] (see e.g. [57, Corollary 4.35]). We consider only so-called uniform families of circuits, i.e., their description does not change significantly for different input sizes [10, 57].
Vector-RAM
Following the definition of [7], the Vector-RAM model is a RAM model with a vector processor that executes operations on vectors of length $n$ (the length can vary). Each vector element/word has bit-width $\mathcal{O}(\log(n))$. Vector operations/instructions are typically assumed to admit circuits with $\mathcal{O}(\log(n))$ depth and polynomial size, i.e., NC[1] circuits. Each vector operation is assumed to take constant time, or “one step”. The vector processor can read/write $n$ consecutive words from/to the main memory, also in constant time. A strong aspect of the model is that Vector-RAM algorithms can be efficiently ported to other parallel models, such as PRAM.
TCU
The $(s^{2},l)$-TCU model [14, 17], is a RAM model equipped with a uniform family of circuits that can multiply two matrices $\bm{\mathrm{A}}$ and $\bm{\mathrm{B}}$ with sizes $n\times s$ and $s\times s$, respectively ($n\geq s)$. The computational cost (time complexity) of invoking the circuit is $\mathcal{O}(ns+l)$, where the first term $ns$ is accounted for loading the data from the main memory and writing the result, while $l$ is a system-dependent parameter that corresponds to the latency of invoking the tensor core. The tensor core can read/write $ns$ consecutive words from/to the main memory in constant time. Each word (and matrix element) consists of $B$ bits, where $B\in O(\log(n))$.
@@ 3 MMV-RAM: a model for matrix multiplication AI accelerators
We propose the MMV-RAM model by judiciously extending the Vector-RAM model with a MMU. As illustrated in Figure 3.1, MMV-RAM contains the following three processing units:
@@
The positive integer $s$ is the only parameter of the model, which corresponds to the (right) matrix multiplication size and it is independent of the input length111 Typical values of $s$ in existing architectures are $16,32,128$ or $256$.. To be more precise, the MMU is a uniform family of circuits  [10, 57] that is assumed to perform a single-step matrix multiplication. The VCU in MMV-RAM implements a fixed set of instructions defined by a uniform family of AC[0] circuits, i.e., boolean circuits with unbounded fan-in, constant depth, and polynomial size (a detailed analysis of these instructions is provided in Section 3.3). Table 3.1 summarizes how existing AI accelerators map on the proposed MMV-RAM model.
@@
NPU (AMD) [27] Systolic array of Compute Tiles Single compute tiles
@@
GPU (AMD) [1] CDNA Matrix Core (MI300 series) SIMD Units
@@
TPU (Google) [29] Matrix Multiply Unit (MXU) Vector Processing Unit (VPU)
Table 3.2: Existing upper and lower bounds for the size of constant-depth, unbounded fan-in circuits for the product of two $s\times s$ matrices.
Gates Depth Size upper bound Size lower bound Input Notes
[…]
{,,$\{\wedge,\vee,\neg,\text{Th}\}$ $\mathcal{O}(d),d\geq 1$ $\widetilde{\mathcal{O}}(ds^{a+\beta\gamma^{d}})$ [45] $\Omega\left(s^{2}\tfrac{\lambda_{d}(s^{2})}{d^{2}}\right)$ [48] $\mathcal{O}(\log(s))$-bit ints $a=\log_{2}(7),\beta\approx 1.6,\gamma\approx 0.5$
@@ 3.1 Work/depth cost model
An MMV-RAM algorithm is evaluated using the work and depth analysis as in the Vector-RAM model [8, 7]. At each step, the machine can execute a scalar instruction, a vector instruction, or a matrix multiplication. To keep the cost model of MMV-RAM simple, we focus on bounding the number of VCU and MMU calls (i.e. step complexity) and their length (element complexity [7]).
The work complexity in each step represents the total number of basic (boolean) operations that are executed, potentially in parallel. In a vector instruction, for example, the work depends both on the length of the input vector and also on the number of edges (size) of the corresponding circuit that implements it. Respectively, in a matrix multiplication operation, the work depends on the number of rows of the (left) input matrix, and the bit-width of the matrix elements.
We assign the following costs for operations:
@@
  • $\mathcal{M}(n)$ is the work for matrix multiplications between two matrices with sizes $n\times s$ and $s\times s$. It scales linearly in $n$, which means that we can write $\mathcal{M}(an)=a\mathcal{M}(n)$ and $\mathcal{M}(n+a)=\mathcal{M}(n)+\mathcal{M}(a)$.
  • $\mathcal{V}(n)$ is the work for vector instructions on vectors of length $n$. The only instructions that are assumed “by default” in the model are $\mathsf{AND}$, $\mathsf{OR}$, and $\mathsf{NOT}$, which correspond to vector variants of ,$\wedge$,$\vee$,∧ , ∨ , and $\neg$, respectively.
An example MMU implementation with the aforementioned properties can be achieved by a TC[0] circuit that implements the classical dot-product based matrix multiplication algorithm. The size of such a circuit is upper bounded by $\mathcal{M}(n)\in O(ns^{2})$. We refer to [45] for details, and to Table 3.2 for an overview of existing upper and lower bounds. If complex vector instructions are used by an algorithm, their associated circuit sizes, depths, fan-in, and fan-out must be explicitly defined, since they affect the total work of an algorithm. In Table 3.3, for example, we describe in detail the VCU instructions that we use for our algorithms, and their corresponding costs.
@@
Memory accesses are modeled as follows. The VCU can read $n$ consecutive words from the main memory in one step, where $n$ is arbitrary. The MMU can also read $ns$ consecutive words from the main memory in a single step. These words are either stored in the $n\times s$ or in the $s\times s$ buffer of the unit. The memory layout of the input and output matrices in MMV-RAM can be any fixed matrix layout, say row or column major. Without loss of generality, we assume that the input matrices are read in row-major order, and the output matrix is also returned in row-major order. Matrix transposition of a matrix having $s$ columns can be implemented as a dedicated vector instruction, if required. Last but not least, irregular and strided memory accesses are not explicitly allowed, but could be implemented with a dedicated AC[0] circuit family in the VCU. In this case, the precise circuit definition must be explicitly mentioned.
@@
Instructions that involve arithmetic operations can increase the number of bits required to store the (intermediate) results. For example, integer vector addition/subtraction can increase the number of bits by one, while integer multiplication can double the number of bits. For each algorithm we report upper bounds on the number of required bits to avoid overflow in the results, and the implications on the total complexity.
Table 3.3: Circuit sizes and fan-in for vector operations of length $n$ used in the main Algorithm 1. Integers (int) have bit-width $B$. All circuits have constant depth, and $s$ is the fixed model parameter. By default, only instructions 1. and 2. are assumed by the MMV-RAM model. The remaining instructions (except $\mathsf{REVSPEC}$) are also typically present in existing architectures, potentially with small variations.
Operation Size Fan-in / -out Description
[…]
9. $\bm{\mathrm{z}}:\text{vec[int]}\leftarrow\mathsf{REVSPEC}(\overline{\bm{\mathrm{x}}}:\text{vec[int]},\overline{\bm{\mathrm{f}}}:\text{vec[int] };s)$ $\mathcal{O}\left(nB(s^{2}+\frac{B}{s})\right)$ $\mathcal{O}(s+B)$ / $\mathcal{O}(sB)$ Reverts misspeculation (Lemma 4.1).
@@ 3.2 The balance between the MMU and VCU
We now provide some arguments that both units of MMV-RAM are necessary to design efficient algorithms: the MMU cannot be efficiently simulated by the VCU, but the VCU provide crucial capabilities that are hard for the MMU. We first recall the foundational circuit complexity lower bound that parity is not in AC[0] [21, 62, 22].
@@
where $c_{1}$ and $c_{2}$ are constants. This allows us to prove a clear separation between the computational power of the MMU and the VCU in the MMV-RAM model. A simple example is the following.
@@
Let $\bm{\mathrm{A}}$ and $\bm{\mathrm{B}}$ be two $B$-bit integer matrices with sizes $n\times s$ and $s\times s$, respectively. Let $\mathcal{A}$ be an algorithm in the MMV-RAM model that uses only the scalar and the vector units to compute the matrix product $\bm{\mathrm{A}}\bm{\mathrm{B}}$. If the work of $\mathcal{A}$ is upper bounded by $s^{O(1)}$, then $\mathcal{A}$ necessarily executes $\Omega(\frac{\log(s)}{\log(\log(s))})$ steps.
@@
If we want to compute the parity of an $s$-bit sequence, we can place the $s$ bits as the first row of $\bm{\mathrm{A}}$, and set $\bm{\mathrm{B}}$ to be the all-ones matrix. The least significant bit of the top-left element of the product $\bm{\mathrm{A}}\bm{\mathrm{B}}$ gives the parity. Since $\mathcal{A}$ only uses the scalar and vector units, it forms an unbounded fan-in Boolean circuit, and therefore Equation 3 gives the result. ∎
While the MMU is evidently stronger than the VCU for certain problems, we can show that it is not strong enough by itself to perform arbitrary calculations. That is, even if the VCU implements a “weaker” class of circuits, it still provides crucial capabilities to the model. One example is the following.
@@
Let $s$ be the MMV-RAM model parameter, and let $m$, $n$ be two positive integers such that $m=ns$. There exists a vector operation which can be efficiently implemented in the VCU, i.e., with an AC[0] circuit of size $m$ and depth $\mathcal{O}(1)$, but it requires $\Omega(n)$ calls to the MMU.

### v2

@@ 3 MMV-RAM: A model for matrix multiplication AI accelerators
In this section we propose the MMV-RAM model by judiciously extending the Vector-RAM model with an MMU. As illustrated in Figure˜3.1, MMV-RAM contains the following three processing units:
@@
The positive integer $s$ is the only parameter of the model, which corresponds to the (right) matrix multiplication size and it is independent of the input length [^1] Typical values of $s$ in existing architectures are $16,32,128$ or $256$.. To be more precise, the MMU is a uniform family of circuits  [10, 61] that is assumed to perform a single-step matrix multiplication. The VCU in MMV-RAM implements a fixed set of instructions defined by a uniform family of AC[0] circuits, i.e., boolean circuits with unbounded fan-in, constant depth, and polynomial size. Table˜3.1 summarizes how existing AI accelerators map on the proposed MMV-RAM model. In Appendix 3.3 we discuss further extensions and variants of MMV-RAM.
@@
NPU (AMD) [30] Systolic array Compute Tile
@@
GPU (AMD) [1] CDNA Matrix Core SIMD Unit
@@
TPU (Google) [32] Matrix Multiply Unit Vector Proc. Unit
@@ 3.1 Work/depth cost model
An MMV-RAM algorithm is evaluated using the work and depth analysis as in the Vector-RAM model [8, 7]. At each step, the machine can execute a scalar instruction, a vector instruction, or a matrix multiplication. To keep the cost model of MMV-RAM simple, we focus on bounding the number of VCU and MMU calls (i.e. step complexity), and their length, which implies the work complexity. The latter represents the total number of (boolean) operations that are executed in each step, potentially in parallel. More formally, the following costs are assigned:
@@
  • $\mathcal{M}(n)$ is the work for matrix multiplications between two matrices with sizes $n\times s$ and $s\times s$.
  • $\mathcal{V}(n)$ is the work for vector instructions on vectors of length $n$.
$\mathcal{M}(n)$ depends on the number of rows of the (left) input matrix, and the bit-width of the matrix elements. It scales linearly in $n$, which means that we can write $\mathcal{M}(an)=a\mathcal{M}(n)$ and $\mathcal{M}(n+a)=\mathcal{M}(n)+\mathcal{M}(a)$. Note that we do not assume a specific circuit implementation for the MMU, but rather define it only based on the aforementioned properties. For example, a constant-depth threshold circuit (TC[0]) that implements the standard dot-product based algorithm satisfies these properties. We refer to [49] and Appendix A.1 for more details on matrix multiplication circuits.
$\mathcal{V}(n)$ depends on the length of the input vector and on the number of edges (size) of the corresponding AC[0] circuit that implements it. To be precise, VCU instructions are uniform families of unbounded fan-in circuits with constant depth and polynomial size in all the involved parameters, i.e., the input length $n$, the number of “bits-per-element” $B$, and the model parameter $s$. The only instructions that are assumed “by default” in the model are element-wise $\wedge,\vee,$ and $\neg$. If complex vector instructions are used by an algorithm, their associated circuit sizes, depths, fan-in, and fan-out must be explicitly defined. Some examples can be found in Appendix A.2.2.
@@
Memory accesses are modeled as follows. The VCU can read $n$ consecutive words from the main memory in one step, where $n$ is arbitrary. The MMU can also read $ns$ consecutive words from the main memory in a single step. These words are either stored in the $n\times s$ or in the $s\times s$ buffer of the unit. The memory layout of the input and output matrices in MMV-RAM can be any fixed matrix layout, say row or column major. Without loss of generality, we assume that the input matrices are read in row-major order, and the output matrix is also returned in row-major order. Matrix transposition of a matrix having $s$ columns can be implemented as a dedicated vector instruction, if required. Last but not least, irregular and strided memory accesses are not explicitly allowed, but they can be implemented with a dedicated AC[0] circuit family in the VCU. In this case, a precise circuit definition is necessary.
@@
Instructions that involve arithmetic operations can increase the number of bits required to store the (intermediate) results. For example, integer addition can increase the number of bits by one, while multiplication can double the number of bits. For each algorithm we upper bound the number of required bits to avoid overflow, since it affects the total complexity.
@@ 3.2 The balance between the MMU and VCU
We now argue that both units of MMV-RAM are necessary to design efficient algorithms: the MMU cannot be efficiently simulated by the VCU, but the VCU provides crucial capabilities that are hard (or even impossible) for the MMU. The foundational result that our analysis builds upon is the circuit complexity lower bound that parity is not in AC[0] [22, 66, 24]. We first recall the following classic theorem by Håstad.
@@
where $c_{1}$ and $c_{2}$ are constants. This allows us to prove a clear separation between the computational power of the MMU and the VCU in the MMV-RAM model. A straightforward example is the following.
@@
Let $\bm{\mathrm{A}}$ and $\bm{\mathrm{B}}$ be two $B$-bit integer matrices with sizes $n\times s$ and $s\times s$, respectively. Let $\mathcal{A}$ be an algorithm in the MMV-RAM model that uses only the scalar and the vector units to compute the matrix product $\bm{\mathrm{A}}\bm{\mathrm{B}}$. If the work of $\mathcal{A}$ is upper bounded by $s^{O(1)}$, then $\mathcal{A}$ executes $\Omega\left(\frac{\log(s)}{\log(\log(s))}\right)$ steps.
@@
If we want to compute the parity of an $s$-bit sequence, we can place the $s$ bits as the first row of $\bm{\mathrm{A}}$, and set $\bm{\mathrm{B}}$ to be the all-ones matrix. The least significant bit of the top-left element of the product $\bm{\mathrm{A}}\bm{\mathrm{B}}$ gives the parity. Since $\mathcal{A}$ only uses the scalar and vector units, it forms an unbounded fan-in Boolean circuit, and therefore Eq. (2) gives the result. ∎
Next, we argue that the MMU is not strong enough by itself, and that the VCU provides crucial capabilities to the model. The first observation is that the MMU can only execute linear transformations. Therefore, it is unable to execute important non-linear functions, such as ReLU. The MMU might be able to approximate such functions, but this is not covered in this work, since ReLU, for example, can easily be implemented with the VCU.
Besides non-linear functions, we can also show that the VCU is important even for linear transformations, even if it implements a “weaker” class of circuits than the MMU. One example is detailed in Lemma 3.4 below.
@@
Let $s$ be the MMV-RAM model parameter, and let $m$, $n$ be two positive integers such that $n=ms$, and $m\gg s$. There exists a vector operation which can be efficiently implemented in the VCU, i.e., with an AC[0] circuit of size $n$ and depth $\mathcal{O}(1)$, but it requires $\Omega(s)$ calls to the MMU.
@@
In the context of this lemma, only two types of operations are allowed: reading and writing consecutive positions between the memory and the MMU, and executing left and right matrix multiplications. In order to model them algebraically, we first introduce some notation.
We always treat the vector $\bm{\mathrm{x}}$, which has length $n$ initially, as the contents of the memory. Using a single read operation, we can load any $ks$ consecutive elements of $\bm{\mathrm{x}}$ in the left “matrix register” of the MMU, for some $k$, starting from some arbitrary position $r$, or $s^{2}$ consecutive positions in the right register of the MMU. In the former case, the $ks$ consecutive elements of $\bm{\mathrm{x}}$ are viewed as a row-major matrix of size $k\times s$. We now want to multiply $\bm{\mathrm{X}}$ with some arbitrary $s\times s$ matrix $\bm{\mathrm{B}}_{s}$ from the right. By using the mixed-product property of the Kronecker product $\otimes$, this operation can be written as follows:
Right multiplication: $\bm{\mathrm{x}}^{\prime}\leftarrow\begin{pmatrix}\bm{\mathrm{I}}_{r-1}&&\\ &\bm{\mathrm{I}}_{k}\otimes\bm{\mathrm{B}}_{s}^{\top}&\\ &&\bm{\mathrm{I}}_{n-ks-r+1}\end{pmatrix}\bm{\mathrm{x}}.$
In a similar way we model the left multiplications. In this case we read $s^{2}$ consecutive elements of $\bm{\mathrm{x}}$, and store them as a row-major $s\times s$ matrix $\bm{\mathrm{X}}$ in the right register of the MMU, starting again at some arbitrary position $r$. $\bm{\mathrm{X}}$ needs to be multiplied from the left with some other matrix $\bm{\mathrm{A}}_{k\times s}$ with size $k\times s$, and then the resulting matrix $\bm{\mathrm{A}}_{k\times s}\bm{\mathrm{X}}$ is stored back in memory. Using again properties of Kronecker products, we can write:
Left multiplication: $\bm{\mathrm{x}}^{\prime}\leftarrow\begin{pmatrix}\bm{\mathrm{I}}_{r-1}&&\\ &\bm{\mathrm{A}}_{k\times s}\otimes\bm{\mathrm{I}}_{s}&\\ &&\bm{\mathrm{I}}_{n-s^{2}-r+1}\end{pmatrix}\bm{\mathrm{x}},$